### v3

@@ 3.1.2 Data
Pretraining data mainly serve two purposes, i.e. (1) aligning different modalities and (2) providing world knowledge. The pretraining corpora can be divided into coarse-grained and fine-grained data according to granularities, which we will introduce sequentially. We summarize commonly used pretraining datasets in Table 4.
@@
SBU Captions [67]. It is a captioned photo dataset containing 1M image-text pairs, with images and descriptions sourced from Flickr. Specifically, an initial set of images is acquired by querying the Flickr website with a large number of query terms. The descriptions attached to the images thus serve as captions. Then, to ensure that descriptions are relevant to the images, the retained images fulfill these requirements: (1) Descriptions of the images are of satisfactory length, decided by observation. (2) Captions should contain at least 2 words in the predefined term lists and a propositional word (e.g. “on”, “under”) that suggests spatial relationships.
@@
  • LAION-COCO [69]: It contains 600M images extracted from the English subset of LAION-5B. The captions are synthetic, using BLIP [70] to generate various image captions and using CLIP [12] to pick the best fit.
Table 4: Common datasets used for pre-training.
@@
COYO-700M [71]. It contains 747M image-text pairs, which are extracted from CommonCrawl. In terms of data filtering, the authors design the following strategies to filter out data samples: (1) For images, those with inappropriate size, content, format, or aspect ratio are filtered. Moreover, the images are filtered based on the pHash value to remove images overlapped with public datasets such as ImageNet and MS-COCO. (2) For text, only English text with satisfactory length, noun forms, and appropriate words are saved. Whitespace before and after the sentence will be removed, and consecutive whitespace characters will be replaced with a single whitespace. Moreover, text appearing more than 10 times (e.g. “image for”) will be dropped. (3) For image-text pairs, duplicated samples are removed based on (image pHash, text) tuple.
Recently, more works [64, 73] have explored generating high-quality fine-grained data through prompting strong MLLMs (e.g. GPT-4V). Compared with coarse-grained data, these data generally contain longer and more accurate descriptions of the images, thus enabling finer-grained alignment between image and text modalities. However, since the approach generally requires calling commercial-use MLLMs, the cost is higher, and the data volume is smaller. Notably, ShareGPT4V [64] strikes a balance by first training a captioner with GPT-4V-generated 100K data, then scaling up the data volume to 1.2M using the pre-trained captioner.
@@ 3.2 Instruction-tuning
Figure 3: Comparison of three typical learning paradigms, adapted from [76].
@@ 3.2.1 Introduction
Instruction refers to the description of tasks. Intuitively, instruction tuning aims to teach models to better understand the instructions from users and fulfill the demanded tasks. Tuning in this way, LLMs can generalize to unseen tasks by following new instructions, thus boosting zero-shot performance. This simple yet effective idea has sparked the success of subsequent NLP works, such as ChatGPT [77], InstructGPT [78].
The comparisons between instruction tuning and related typical learning paradigms are illustrated in Fig. 3. The supervised fine-tuning approach usually requires a large amount of task-specific data to train a task-specific model. The prompting approach reduces the reliance on large-scale data and can fulfill a specialized task via prompt engineering. In such a case, though the few-shot performance has been improved, the zero-shot performance is still quite average [4]. Differently, instruction tuning learns how to generalize to unseen tasks rather than fitting specific tasks like the two counterparts. Moreover, instruction tuning is highly related to multi-task prompting [79] and learning [80].
@@ 3.2.2 Training detail
Table 5: A simplified template to structure the multimodal instruction data. <instruction> is a textual description of the task. {<image>, <text>} and <output> are input and output from the data sample. Note that <text> in the input may be missed for some datasets, such as image caption datasets merely have <image>. The example is adapted from [81].
A multimodal instruction sample often includes an optional instruction and an input-output pair. The instruction is typically a natural language sentence describing the task, such as, “Describe the image in detail.” The input can be an image-text pair like the VQA task [82] or only an image like the image caption task [83]. The output is the answer to the instruction conditioned on the input. The instruction template is flexible and subject to manual designs [21], as exemplified in Table 5. Note that the instruction template can also be generalized to the case of multi-round human-agent conversations [81, 16].
@@
Here, $\mathcal{A}$ denotes the predicted answer, and $\theta$ are the parameters of the model. The training objective is typically the original auto-regressive objective used to train LLMs [16], based on which the MLLM is encouraged to predict the next token of the response sequentially:
@@
where $N$ is the length of the ground-truth.
@@ 3.2.3 Data collection
Table 6: Instruction templates for VQA datasets, cited from [51]. <Image> and {Question} are the image and the question in the original VQA datasets, respectively.
Data Adaptation. Task-specific datasets are rich sources of high-quality data. Hence, abundant works [51, 84] have utilized existing high-quality datasets to construct instruction-formatted datasets. Take the transformation of VQA datasets for an example: the original sample is an input-out pair where the input comprises an image and a natural language question, and the output is the textual answer to the question conditioned on the image. The input-output pairs of these datasets could naturally comprise the multimodal input and response of the instruction sample. The instructions, i.e. the descriptions of the tasks, can either derive from manual design or from semi-automatic generation aided by GPT. Specifically, some works [17] hand-craft a pool of candidate instructions and sample one of them during training. We offer an example of instruction templates for the VQA datasets as shown in Table 6. The other works manually design some seed instructions and use these to prompt GPT to generate more [21].
Note that since the answers of existing VQA and caption datasets are usually concise, directly using these datasets for instruction tuning may limit the output length of MLLMs. There are two common strategies to tackle this problem. The first one is to specify the corresponding requirements explicitly in the instructions. For example, ChatBridge [85] explicitly declares short and brief for short-answer data. The second one is to extend the length of existing answers [86]. For example, M3IT [86] proposes to rephrase the original answer by prompting ChatGPT with the original question, answer, and contextual information of the image (e.g. caption and text extracted through OCR).
Table 7: A summary of popular datasets generated by self-instruction. For input/output modalities, I: Image, T: Text, V: Video, A: Audio. For data composition, M-T and S-T denote multi-turn and single-turn, respectively.
@@
LLaVA-Instruct [16] 158K I + T $\rightarrow$ T MS-COCO 23K caption + 58K M-T QA + 77K reasoning
LVIS-Instruct [72] 220K I + T $\rightarrow$ T LVIS 110K caption + 110K M-T QA
ALLaVA [73] 1.4M I + T $\rightarrow$ T VFlan, LAION 709K caption + 709K S-T QA
Video-ChatGPT [87] 100K V + T $\rightarrow$ T ActivityNet 7K description + 4K M-T QA
VideoChat [21] 11K V+T $\rightarrow$ T WebVid description + summarization + creation
Clotho-Detail [88] 3.9K A + T $\rightarrow$ T Clotho caption
Self-Instruction. Although existing multi-task datasets can contribute a rich source of data, they usually do not meet human needs well in real-world scenarios, such as multiple-round conversations. To tackle this issue, some works collect samples through self-instruction [89], which utilizes LLMs to generate textual instruction-following data using a few hand-annotated samples. Specifically, some instruction-following samples are hand-crafted as demonstrations, after which ChatGPT/GPT-4 is prompted to generate more instruction samples with the demonstrations as guidance. LLaVA [16] extends the approach to the multimodal field by translating images into text of captions and bounding boxes, and prompting text-only GPT-4 to generate new data with the guidance of requirements and demonstrations. In this way, a multimodal instruction dataset is constructed, called LLaVA-Instruct-150k. Following this idea, subsequent works such as MiniGPT-4 [17] and GPT4Tools [90] develop different datasets catering to different needs. Recently, with the release of the more powerful multimodal model GPT-4V, many works have adopted GPT-4V to generate data of higher quality, as exemplified by LVIS-Instruct4V [72] and ALLaVA [73]. We summarize the popular datasets generated through self-instruction in Table 7. It should be noted that this paradigm highly relies on advanced yet close-sourced models, which can be expensive for data scaling. This approach might be partially due to the limited capabilities of early models. Future research can explore capitalizing on open-sourced models to generate high-quality instruction data.
Data Mixture. Apart from the multimodal instruction data, language-only user-assistant conversation data can also be used to improve conversational proficiencies and instruction-following abilities [91]. LaVIN [91] directly constructs a minibatch by randomly sampling from both language-only and multimodal data. MultiInstruct [84] probes different strategies for training with a fusion of single modal and multimodal data, including mixed instruction tuning (combine both types of data and randomly shuffle) and sequential instruction tuning (text data followed by multimodal data).
@@ 3.2.4 Data quality
Recent research has revealed that the data quality of instruction-tuning samples is no less important than quantity. Lynx [62] finds that models pre-trained on large-scale but noisy image-text pairs do not perform as well as models pre-trained with smaller but cleaner datasets. Similarly, Wei et al.  [92] finds that less instruction-tuning data with higher quality can achieve better performance. For data filtering, the work proposes some metrics to evaluate data quality and, correspondingly, a method to automatically filter out inferior vision-language data. Here we discuss two important aspects of data quality.
@@
Task Coverage. In terms of tasks involved in training data, Du et al.  [93] perform an empirical study and find that the visual reasoning task is superior to captioning and QA tasks for boosting model performance. Moreover, the study suggests that more complex instructions are better than increasing task diversity and incorporating fine-grained spatial annotations.
@@ 3.3.1 Introduction
Alignment tuning is more often used in scenarios where models need to be aligned with specific human preferences, e.g. response with fewer hallucinations. Currently, Reinforcement Learning with Human Feedback (RLHF) and Direct Preference Optimization (DPO) are two main techniques for alignment tuning. In this section, we introduce the main ideas of the two techniques in sequence and offer some examples of how they are utilized in addressing practical problems, and finally, give a compilation of the related datasets.
@@ 3.3.2 Training detail
  1. (1)
    Supervised fine-tuning. This step aims to fine-tune a pre-trained model to present the preliminary desired output behavior. The fine-tuned model in the RLHF setting is called a policy model. Note that this step might be skipped since the supervised policy model $\pi^{\text{SFT}}$ can be initialized from an instruction-tuned model.
  2. (2)
    Reward modeling. A reward model is trained using preference pairs in this step. Given a multimodal prompt (e.g. image and text) $x$ and a response pair $(y_{w},y_{l})$, the reward model $r_{\theta}$ learns to give a higher reward to the preferred response $y_{w}$, and vice versa for $y_{l}$, with the objective:
    $\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}&\Big{[}\text{log}(\sigma(r_{\theta}(x,y_{w})\\ &-r_{\theta}(x,y_{l})\Big{]}\end{split}$ (3)
@@
    $\mathcal{L}(\phi)$ $=-\mathbb{E}_{x\sim\mathcal{D},y\sim\pi^{RL}_{\phi}(y|x)}\Big{[}r_{\theta}(x,y)$ (4)
    $-\beta\cdot\mathbb{D}_{KL}\Big{(}\pi^{RL}_{\phi}(y|x)||\pi^{REF}(y|x)\Big{)}\Big{]}$
@@
DPO [97]. It learns from human preference labels utilizing a simple binary classification loss. Compared with the PPO-based RLHF algorithm, DPO is exempt from learning an explicit reward model, thus simplifying the whole pipeline to two steps, i.e. human preference data collection and preference learning. The learning objective for the algorithm is as follows:
@@ 3.3.3 Data
The gist of data collection for alignment-tuning is to collect feedback for model responses, i.e. to decide which response is better. It is generally more expensive to collect such data, and the amount of data used for this phase is typically even less than that used in previous stages. In this part, we introduce some datasets and summarize them in Table 8.
LLaVA-RLHF [96]. It contains 10K preference pairs collected from human feedback in terms of honesty and helpfulness. The dataset mainly serves to reduce hallucinations.
RLHF-V [98]. It has 5.7K fine-grained human feedback data collected by performing segment-level hallucination corrections.
@@
Table 8: A summary of datasets for alignment-tuning. For input/output modalities, I: Image, T: Text.
@@ 4 Evaluation
Evaluation is an essential part of developing MLLMs since it provides feedback for model optimization and helps to compare the performance of different models. Compared with evaluation methods of traditional multimodal models, the evaluation of MLLMs exhibits several new traits: (1) Since MLLMs are generally versatile, it is important to evaluate MLLMs comprehensively. (2) MLLMs exhibit many emergent capabilities that require special attention (e.g. OCR-free math reasoning) and thus require new evaluation schemes. The evaluation of MLLMs can be broadly categorized into two types according to the question genres, including closed-set and open-set. Closed-set evaluation often involves task-specific benchmarks and more comprehensive benchmarks specifically designed for MLLM, where answers are limited to predefined sets. Open-set evaluation typically includes manual scoring, GPT scoring, and case study.

### v4

@@ 3.1.2 Data
Pretraining data mainly serve two purposes, i.e. (1) aligning different modalities and (2) providing world knowledge. The pretraining corpora can be divided into coarse-grained and fine-grained data according to granularities, which we will introduce sequentially. We summarize commonly used pretraining datasets in Table IV.
@@
SBU Captions [86]. It is a captioned photo dataset containing 1M image-text pairs, with images and descriptions sourced from Flickr. Specifically, an initial set of images is acquired by querying the Flickr website with a large number of query terms. The descriptions attached to the images thus serve as captions. Then, to ensure that descriptions are relevant to the images, the retained images fulfill these requirements: (1) Descriptions of the images are of satisfactory length, decided by observation. (2) Descriptions of the images contain at least 2 words in the predefined term lists and a propositional word (e.g. “on”, “under”) that generally suggests spatial relationships.
@@
  • LAION-COCO [88]: It contains 600M images extracted from the English subset of LAION-5B. The captions are synthetic, using BLIP [89] to generate various image captions and using CLIP [13] to pick the best fit for the image.
TABLE IV: Common datasets used for pre-training.
@@
COYO-700M [90]. It contains 747M image-text pairs, which are extracted from CommonCrawl. For data filtering, the authors design the following strategies: (1) For images, those with inappropriate size, content, format, or aspect ratio are filtered. Moreover, the images are filtered based on the pHash value to remove images overlapped with public datasets such as ImageNet and MS-COCO. (2) For text, only English text with satisfactory length, noun forms, and appropriate words are saved. Whitespace before and after the sentence will be removed, and consecutive whitespace characters will be replaced with a single whitespace. Moreover, text appearing more than 10 times (e.g. “image for”) will be dropped. (3) For image-text pairs, duplicated samples are removed based on (image pHash, text) tuple.
Recently, more works [83, 91, 92] have explored generating high-quality fine-grained data through prompting strong MLLMs (e.g. GPT-4V). Compared with coarse-grained data, these data generally contain longer and more accurate descriptions of the images, thus enabling finer-grained alignment between image and text modalities. However, since this approach generally requires calling commercial-use MLLMs, the cost is higher, and the data volume is relatively smaller. Notably, ShareGPT4V [83] strikes a balance by first training a captioner with GPT-4V-generated 100K data, then scaling up the data volume to 1.2M using the pre-trained captioner.
@@ 3.2 Instruction-tuning
Figure 3: Comparison of three typical learning paradigms. The image is from [19].
@@ 3.2.1 Introduction
Instruction refers to the description of tasks. Intuitively, instruction tuning aims to teach models to better understand the instructions from users and fulfill the demanded tasks. Tuning in this way, LLMs can generalize to unseen tasks by following new instructions, thus boosting zero-shot performance. This simple yet effective idea has sparked the success of subsequent NLP works, such as ChatGPT [2], InstructGPT [95], FLAN [56, 19], and OPT-IML [96].
The comparisons between instruction tuning and related typical learning paradigms are illustrated in Fig. 3. The supervised fine-tuning approach usually requires a large amount of task-specific data to train a task-specific model. The prompting approach reduces the reliance on large-scale data and can fulfill a specialized task via prompt engineering. In such a case, though the few-shot performance has been improved, the zero-shot performance is still quite average [7]. Differently, instruction tuning learns how to generalize to unseen tasks rather than fitting specific tasks like the two counterparts. Moreover, instruction tuning is highly related to multi-task prompting [97].
@@ 3.2.2 Training Detail
TABLE V: A simplified template to structure the multimodal instruction data. <instruction> is a textual description of the task. {<image>, <text>} and <output> are input and output from the data sample. Note that <text> in the input may be missed for some datasets, such as image caption datasets merely have <image>. The example is adapted from [98].
A multimodal instruction sample often includes an optional instruction and an input-output pair. The instruction is typically a natural language sentence describing the task, such as, “Describe the image in detail.” The input can be an image-text pair like the VQA task [99] or only an image like the image caption task [100]. The output is the answer to the instruction conditioned on the input. The instruction template is flexible and subject to manual designs [25, 98, 20], as exemplified in Table V. Note that the instruction template can also be generalized to the case of multi-round conversations [98, 37, 20, 71].
@@
Here, $\mathcal{A}$ denotes the predicted answer, and $\theta$ are the parameters of the model. The training objective is typically the original auto-regressive objective used to train LLMs [20, 37, 71, 101], based on which the MLLM is encouraged to predict the next token of the response. The objective can be expressed as:
@@
where $N$ is the length of the ground-truth response.
@@ 3.2.3 Data Collection
TABLE VI: Instruction templates for VQA datasets, cited from [60]. <Image> and {Question} are the image and the question in the original VQA datasets, respectively.
Data Adaptation. Task-specific datasets are rich sources of high-quality data. Hence, abundant works [60, 82, 70, 102, 76, 103, 104, 101] have utilized existing high-quality datasets to construct instruction-formatted datasets. Take the transformation of VQA datasets for an example, the original sample is an input-out pair where the input comprises an image and a natural language question, and the output is the textual answer to the question conditioned on the image. The input-output pairs of these datasets could naturally comprise the multimodal input and response of the instruction sample (see §3.2.2). The instructions, i.e. the descriptions of the tasks, can either derive from manual design or from semi-automatic generation aided by GPT. Specifically, some works [60, 102, 21, 70, 35, 105] hand-craft a pool of candidate instructions and sample one of them during training. We offer an example of instruction templates for the VQA datasets as shown in Table VI. The other works manually design some seed instructions and use these to prompt GPT to generate more [82, 25, 98].
Note that since the answers of existing VQA and caption datasets are usually concise, directly using these datasets for instruction tuning may limit the output length of MLLMs. There are two common strategies to tackle this problem. The first one is to specify explicitly in instructions. For example, ChatBridge [104] explicitly declares short and brief for short-answer data, as well as a sentence and single sentence for conventional coarse-grained caption data. The second one is to extend the length of existing answers [105]. For example, M3IT [105] proposes to rephrase the original answer by prompting ChatGPT with the original question, answer, and contextual information of the image (e.g. caption and OCR).
TABLE VII: A summary of popular datasets generated by self-instruction. For input/output modalities, I: Image, T: Text, V: Video, A: Audio. For data composition, M-T and S-T denote multi-turn and single-turn, respectively.
@@
LLaVA-Instruct 158K I + T $\rightarrow$ T MS-COCO 23K caption + 58K M-T QA + 77K reasoning
LVIS-Instruct 220K I + T $\rightarrow$ T LVIS 110K caption + 110K M-T QA
ALLaVA 1.4M I + T $\rightarrow$ T VFlan, LAION 709K caption + 709K S-T QA
Video-ChatGPT 100K V + T $\rightarrow$ T ActivityNet 7K description + 4K M-T QA
VideoChat 11K V+T $\rightarrow$ T WebVid description + summarization + creation
Clotho-Detail 3.9K A + T $\rightarrow$ T Clotho caption
Self-Instruction. Although existing multi-task datasets can contribute a rich source of data, they usually do not meet human needs well in real-world scenarios, such as multiple rounds of conversations. To tackle this issue, some works collect samples through self-instruction [106], which utilizes LLMs to generate textual instruction-following data using a few hand-annotated samples. Specifically, some instruction-following samples are hand-crafted as demonstrations, after which ChatGPT/GPT-4 is prompted to generate more instruction samples with the demonstrations as guidance. LLaVA [20] extends the approach to the multimodal field by translating images into text of captions and bounding boxes, and prompting text-only GPT-4 to generate new data with the guidance of requirements and demonstrations. In this way, a multimodal instruction dataset is constructed, called LLaVA-Instruct-150k. Following this idea, subsequent works such as MiniGPT-4 [21], ChatBridge [104], GPT4Tools [107], and DetGPT [72] develop different datasets catering for different needs. Recently, with the release of the more powerful multimodal model GPT-4V, many works have adopted GPT-4V to generate data of higher quality, as exemplified by LVIS-Instruct4V [91] and ALLaVA [92]. We summarize the popular datasets generated through self-instruction in Table VII.
Data Mixture. Apart from the multimodal instruction data, language-only user-assistant conversation data can also be used to improve conversational proficiencies and instruction-following abilities [103, 81, 98, 101]. LaVIN [101] directly constructs a minibatch by randomly sampling from both language-only and multimodal data. MultiInstruct [102] probes different strategies for training with a fusion of single modal and multimodal data, including mixed instruction tuning (combine both types of data and randomly shuffle) and sequential instruction tuning (text data followed by multimodal data).
@@ 3.2.4 Data Quality
Recent research has revealed that the data quality of instruction-tuning samples is no less important than quantity. Lynx [73] finds that models pre-trained on large-scale but noisy image-text pairs do not perform as well as models pre-trained with smaller but cleaner datasets. Similarly, Wei et al. [108] finds that less instruction-tuning data with higher quality can achieve better performance. For data filtering, the work proposes some metrics to evaluate data quality and, correspondingly, a method to automatically filter out inferior vision-language data. Here we discuss two important aspects regarding data quality.
@@
Task Coverage. In terms of tasks involved in training data, Du et al. [109] perform an empirical study and find that the visual reasoning task is superior to captioning and QA tasks for boosting model performance. Moreover, the study suggests that enhancing the complexity of instructions might be more beneficial than increasing task diversity and incorporating fine-grained spatial annotations.
@@ 3.3.1 Introduction
Alignment tuning is more often used in scenarios where models need to be aligned with specific human preferences, e.g. response with fewer hallucinations (see §6). Currently, Reinforcement Learning with Human Feedback (RLHF) and Direct Preference Optimization (DPO) are two main techniques for alignment tuning. In this section, we introduce the main ideas of the two techniques in sequence and offer some examples of how they are utilized in addressing practical problems, and finally, give a compilation of the related datasets.
@@ 3.3.2 Training Detail
  1. 1.
    Supervised fine-tuning. This step aims to fine-tune a pre-trained model to present the preliminary desired output behavior. The fine-tuned model in the RLHF setting is called a policy model. Note that this step might be skipped since the supervised policy model $\pi^{\text{SFT}}$ can be initialized from an instruction-tuned model (see §3.2).
  2. 2.
    Reward modeling. A reward model is trained using preference pairs in this step. Given a multimodal prompt (e.g. image and text) $x$ and a response pair $(y_{w},y_{l})$, the reward model $r_{\theta}$ learns to give a higher reward to the preferred response $y_{w}$, and vice versa for $y_{l}$, according to the following objective:
    $\mathcal{L}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log(\sigma(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\right]$ (3)
@@
    $\begin{split}\mathcal{L}(\phi)&=-\mathbb{E}_{x\sim\mathcal{D},y\sim\pi^{RL}_{\phi}(y|x)}\Big{[}r_{\theta}(x,y)\\ &-\beta\cdot\mathbb{D}_{KL}\Big{(}\pi^{RL}_{\phi}(y|x)||\pi^{REF}(y|x)\Big{)}\Big{]}\end{split}$ (4)
@@
DPO [113]. It learns from human preference labels utilizing a simple binary classification loss. Compared with the PPO-based RLHF algorithm, DPO is exempt from learning an explicit reward model, thus simplifying the whole pipeline to two steps, i.e. human preference data collection and preference learning. The learning objective is as follows:
@@ 3.3.3 Data
The gist of data collection for alignment-tuning is to collect feedback for model responses, i.e. to decide which response is better. It is generally more expensive to collect such data, and the amount of data used for this phase is typically even less than that used in previous stages. In this part, we introduce some datasets and summarize them in Table VIII.
LLaVA-RLHF [112]. It contains 10K preference pairs collected from human feedback in terms of honesty and helpfulness. The dataset mainly serves to reduce hallucinations in model responses.
RLHF-V [114]. It has 5.7K fine-grained human feedback data collected by segment-level hallucination corrections.
@@
TABLE VIII: A summary of datasets for alignment-tuning. For input/output modalities, I: Image, T: Text.
@@ 4 Evaluation
Evaluation is an essential part of developing MLLMs since it provides feedback for model optimization and helps to compare the performance of different models. Compared with evaluation methods of traditional multimodal models, the evaluation of MLLMs exhibits several new traits: (1) Since MLLMs are generally versatile, it is important to evaluate MLLMs comprehensively. (2) MLLMs exhibit many emergent capabilities that require special attention (e.g. OCR-free math reasoning) and thus require new evaluation schemes. The evaluation of MLLMs can be broadly categorized into two types according to the question genres, including closed-set and open-set.